Calculate NNP at market price and private income from the following;

Net current transfers to abroad& & & & & 30
Mixed income& & & & & & & & & & & & & & & & & & & & & & & & & & & & & 600
Subsidies& & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & 20
Operating surplus& & & & & & & & & & & & & & & & & & & & & & & & & & 200
National debt interest& & & & & & & & & & & & & & & & & & & & 70
Net factor income to abroad& & & & & & & & & & 10
Compensation of employees& & & & & & & & & & 1400
Indirect tax& & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & & 100
Domestic product accruing to govt& 350
Current transfers by govt& & & & & & & & & & & & & & & & 50
1
Expert's answer
2012-03-16T11:25:22-0400
Calculate NNP at market price and private income from the
following;
Net current transfers to abroad 30
Mixed income
600
Subsidies 20
Operating surplus 200
National debt interest 70
Net
factor income to abroad 10
Compensation of employees 1400
Indirect tax
100
Domestic product accruing to govt 350
Current transfers by govt
50
NNP = 600 + 20 + 200 + 70 + 1400 + 100 + 350 - 30 - 50 = 2680
Private
income = 600 + 1400 = 2000